OpenStudy (cookie_2046):

A rectrangle has a length of n units. It's width is 8 units more than it's length what is the area of the rectangle A. 2 + 8n B. n(n - 8) C. n(n + 8) D. 2n + 8

Vocaloid (vocaloid):

length = n width = n + 8 area = (length)*(width) = ?

OpenStudy (cookie_2046):

c?

Vocaloid (vocaloid):

yeah

OpenStudy (cookie_2046):

Ok Thanks for helping!

OpenStudy (cookie_2046):

Oh and it said that it was correct (: